# M2NO: Multiresolution Operator Learning with Multiwavelet-based Algebraic Multigrid Method

### v2

@@ Appendix E Ablation and Hyperparameter Study
w/o multilevel 4.53E-05
w/o multiwavelet 4.56E-05
M2NO-LPF 7.52E-06
M2NO (Baseline) 7.16E-06
Ablation Study Analysis. In our ablation study of the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO) on Burgers’ equation with $\nu=1e-3$ and a resolution of 1024, summarized in Table 10, we investigated the impact of different model configurations on $L_{2}$ error. The baseline M2NO, utilizing full multilevel and multiwavelet features, achieved an $L_{2}$ error of 7.16E-06. Removal of the multilevel component or the multiwavelet feature resulted in similar increases in error to 4.53E-05 and 4.56E-05, respectively, underscoring their critical roles in reducing error. The M2NO-LPF variant, employing a low-pass filter instead of a traditional multigrid operator, demonstrated a slight increase in error to 7.52E-06, suggesting that while LPF simplifies the model, it retains much of the baseline’s efficacy, particularly in smoothing high-frequency details. These findings highlight the essential contributions of multilevel and multiwavelet components to the M2NO framework and suggest the potential for simpler models in specific scenarios.
Table 11: Hyperparameter study results for the M2NO model, detailing the $L_{2}$ error across various configurations. Changes in grid level, wavelet number, basis number, and layer number are evaluated against the baseline to assess their impact on model performance.
@@
Hyperparameter Study Analysis. The hyperparameter study of M2NO, presented in Table 11, explored the impact of various model configurations on the $L_{2}$ error. Adjustments were made across multiple parameters, including iteration steps, grid levels, wavelet numbers, basis numbers, and the number of layers. Notably, increasing the grid levels from a flat [1,1,1,1,1,1] to a varied [1,2,3] configuration reduced the $L_{2}$ error from 7.35E-06 to 6.45E-06. Changes in the wavelet number revealed that using c=16 minimized the error to 5.29E-06, indicating a more significant wavelet count enhances accuracy. Similarly, an optimal layer count was identified, with six layers achieving the lowest error at 4.66E-06. Conversely, varying the basis number showed that a higher count does not always correlate with lower error, as k=4 performed better than k=2. These results illustrate the nuanced impacts of each parameter on model performance, guiding further refinement and optimization of M2NO configurations for enhanced precision in scientific simulations.
Appendix F Implementation Details
F.1 Baseline and Benchmark Construction
Time-dependent tasks. In our benchmarks, all tasks are time-dependent, except for Darcy Flow. These tasks are auto-regressive, requiring sequential predictions for each subsequent time step. Accordingly, we employ the 2D versions of the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), along with the Fourier Neural Operator (FNO) and Multiwavelet Transform (MWT) models, to predict the next time frame in a step-by-step manner across all benchmarks.
Data splits. For the Diffusion-Reaction and Navier-Stokes (NS) datasets, the data is allocated with 80% used for training, 10% for validation, and 10% for testing. For all other datasets, the split is 60% for training, 20% for validation, and 20% for testing.
Baselines. We implemented all baseline models using their official source code and corresponding default parameters. However, for the LSM model, the default patch size of $[4,4]$ was not suitable for a resolution of $32\times 32$. Consequently, we adjusted the patch size to $[2,2]$ to ensure compatibility. For fairness in comparison, all methods were trained over 500 epochs using the $L_{2}$ loss function [22] and the ADAM optimizer [18] with an initial learning rate of $10^{-3}$, which was halved every 100 epochs ($\gamma=0.5$). Batch sizes were set at 512 for 1D PDEs and 16 for 2D PDEs. All experiments were conducted on a single Nvidia A800 80GB GPU.
F.2 Model Architecture
In the Multiwavelet-based Multigrid Neural Operator (M2NO), the neural networks $A$, $S$, $B$, $C$ and $D^{-1}$ are implemented as single-layer convolutional neural networks (CNNs). The orthonormal polynomial (OP) basis utilized is the Legendre (Leg) basis, chosen with both uniform and non-uniform measures $\mu_{0}$, in alignment with the Multiwavelet Transform (MWT)[12] methodologies.
From a computational perspective, the $restriction$ and $prolongation$ operators depicted in (16) and Fig 1 specifically target the transformation of multiscale and multiwavelet coefficients. Despite this focus on coefficients, the actual input and output of the model are point-wise function samples, specifically $(a_{i},u_{i})$. To address this, we synthesize the data sequence into conceptual functions $f_{a}=\sum_{i=1}^{N}a_{i}\phi_{ji}^{n}$ and $f_{u}=\sum_{i=1}^{N}u_{i}\phi_{ji}^{n}$, positing them within the space $V_{n}^{k}$ where $n=\log_{2}N$. It is critical to note that these functions, $f_{a}$ and $f_{u}$, are employed primarily for theoretical articulation and not for direct computation, serving as a notational convention rather than functional entities in the model’s operations.
Generalizing Dimensionality According to Discrete Wavelet Transform (DWT) theory, the dimension of the analyzed matrix must be a power of 2 [27]. Within the Algebraic Multigrid (AMG) context, such a restriction could severely limit the method’s applicability. Consequently, we generalize as follows: the dimension of the restriction operator $I_{h}^{2h}$, for all $k$, is defined to be $(N/2)\times N$ where
$N=\begin{cases}n,&\text{if }n\text{ is even},\\ n+1,&\text{if }n\text{ is odd},\end{cases}$ (79)
with $n=\text{dim}(A_{j})$. If $n$ is odd, the dimensions of the matrix and the corresponding right-hand side vector are incremented by one.
F.3 Model Configurations
The primary configurations for the Multiwavelet-based Multigrid Neural Operator (M2NO) are detailed in Table 12, and those for the M2NO with Low-Pass Filter (M2NO-LPF) in Table 13. Except where specifically noted, model parameters remain consistent across different hyperparameters and resolutions within the same benchmark.
Table 12: Model Configurations of M2NO.
Modules Advection Burgers Navier-Stokes Darcy Diffusion-Reaction
num_layer 5 5 5 5 5
c 8 8 2 1 4
k 8 8 8 4 4
grid_levels [1,1,2,2,2,3,3] [1,1,1,2,2,2,3] [1,1,2,2,2,3] [3,3,2,2,1] [2,2,1,1,1]
padding_mode ’circular’ ’circular’ ’zeros’ ’zeros’ ’zeros’
bias TRUE TRUE TRUE TRUE TRUE
base ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’
initializer ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’
norm TRUE TRUE TRUE TRUE TRUE
operator ’schur’ ’schur’ ’schur’ ’schur’ ’schur’
activation ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’
Table 13: Model Configurations of M2NO-LPF.
Modules Advection Burgers Navier-Stokes-1e-3 Navier-Stokes-1e-4 Navier-Stokes-1e-5 Darcy Diffusion-Reaction
num_layer 5 5 5 5 5 5 4
c 8 8 1 1 2 2 1
k 8 8 8 8 8 4 8
grid_levels [1,1,1,2,2,2,3] [1,1,1,2,2,2,3] [1,1,2,2,2,3] [1,1,2,2,2,3] [1,1,2,2,2,3] [3,3,2,2,1] [3,3,2,2,2]
padding_mode ’circular’ ’circular’ ’circular’ ’circular’ ’circular’ ’circular’ ’circular’
bias TRUE TRUE TRUE TRUE TRUE TRUE TRUE
base ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’
initializer ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’
norm TRUE TRUE TRUE TRUE TRUE TRUE TRUE
operator ’lpf’ ’lpf’ ’lpf’ ’lpf’ ’lpf’ ’lpf’ ’lpf’
activation ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’

### v3

@@ 4.6. Ablation Study Analysis
w/o Multilevel 4.47E-03
w/o Multigrid 4.55E-03
w/o Multi-wavelet 4.36E-03
FFT as Mapping Operator 3.97E-03
M2NO (Baseline) 3.61E-03
4.7. Hyperparameter Study Analysis
The hyperparameter study of M2NO, presented in Table 7, explored the impact of various model configurations on the $L_{2}$ error. Adjustments were made across multiple parameters, including iteration steps, grid levels, wavelet numbers, basis numbers, and the number of layers. Notably, increasing the grid levels from a flat [1,1,1,1,1,1] to a varied [1,2,3] configuration reduced the $L_{2}$ error from 7.35E-06 to 6.45E-06. Changes in the wavelet number revealed that using c=16 minimized the error to 5.29E-06, indicating a more significant wavelet count enhances accuracy. Similarly, an optimal layer count was identified, with six layers achieving the lowest error at 4.66E-06. Conversely, varying the basis number showed that a higher count does not always correlate with lower error, as k=4 performed better than k=2. These results illustrate the nuanced impacts of each parameter on model performance, guiding further refinement and optimization of M2NO configurations for enhanced precision in scientific simulations.
Table 7. Hyperparameter study results for the M2NO model.
@@ 4.7. Hyperparameter Study Analysis
5. Related Work
Recent advances in neural operators have explored multigrid and wavelet-based architectures to improve efficiency and multi-scale modeling for PDEs. Below, we compare M2NO with representative multigrid and wavelet neural operator approaches.
Multigrid Neural Operators. MgNO (23:mgno) applies multigrid hierarchies to neural operators for efficient linear operator parameterization, but it lacks multiresolution analysis and is limited to linear PDEs. Methods like UGrid (24:UGrid) and Hsieh et al. (19:LearningNP) similarly accelerate linear solvers via multigrid-inspired neural modules with convergence guarantees, yet remain constrained to linear regimes without explicit multi-scale error correction. In contrast, M2NO integrates multiwavelet-based restriction and prolongation within a learnable multigrid cycle, enabling coarse-to-fine error correction for nonlinear, multiscale PDEs.
Wavelet-based Neural Operators. MWT (21:mwt) and its variants, including CMWOL (23:CoupledMO) and WDNO (25:WaveletDNO), leverage multiwavelet transforms to capture localized and multi-scale features, but do not incorporate an explicit hierarchical multigrid structure for iterative error propagation. M2NO advances beyond these by unifying multiwavelet representation with a full multigrid correction framework, supporting efficient error management across all scales.
Technically, M2NO distinguishes itself by combining multiwavelet-based multiresolution analysis with an explicit, learnable multigrid hierarchy. This enables robust coarse-to-fine error correction for nonlinear and multiscale PDEs, bridging gaps left by prior multigrid and wavelet neural operator methods.
6. Conclusion
We introduced M2NO, a multigrid neural operator whose restriction and prolongation steps use predefined multiwavelet filters. This hybrid architecture (i) captures fine–to–coarse interactions in a unified multi-scale representation, (ii) achieves state-of-the-art accuracy with low parameter cost on diverse 1-D/2-D PDE benchmarks (including ERA5), and (iii) can serve as an efficient preconditioner for classical iterative solvers.
Limitation. (i) M2NO currently assumes structured grids; accuracy degrades on highly non-uniform or adaptive meshes. (ii) Very deep multigrid cycles may incur communication overhead on distributed hardware.
Future work. We will (i) extend M2NO to unstructured and adaptive grids by coupling graph-based wavelet kernels with geometric multigrid, (ii) develop dynamic level-of-detail schedules that adjust wavelet depth online, and (iii) systematically investigate its preconditioning power in large-scale scientific simulations. These directions aim to broaden M2NO’s applicability to a wider class of scientific and engineering problems while further improving computational efficiency.
Acknowledgements.
Wei Wang was supported by Advanced Materials-National Science and Technology Major Project (Grant No. 2025ZD0620100), and the Key Research and Development Program in Sichuan Province of China (No. 2024YFFK0410). Zhilu Lai was supported by the Guangdong Provincial Fund - Special Innovation Project (2024KTSCX038) and Research Grants Council of Hong Kong through the Research Impact Fund (R5006-23). Xiaobo Zhang was supported by the Key Research and Development Program in Sichuan Province of China (No. 2024YFFK0410).
Appendix A Multigrid Methods
  • A restriction operator $I_{h}^{2h}$ transfers vectors from the fine grid to the coarse grid.
  • An interpolation operator $I_{2h}^{h}$ returns to the fine grid.
The spectral radius $\rho(I_{h}^{2h})$ is called the convergence factor. It is roughly the worst factor by which the error is reduced with each relaxation sweep.
A.1. Multigrid Theory
Recall the variational properties:
(19) $A^{2h}$ $=I^{2h}_{h}A^{h}I^{h}_{2h}$ $(\text{Galerkin property}),$
$I^{2h}_{h}$ $=c(I^{h}_{2h})^{T},$ $c\in I_{h}^{2h}.$
(20) $\text{Nullspace}(I^{2h}_{h})\;\bot\;\text{Range}[(I^{2h}_{h})^{T}].$
By the second variational property, it then follows that:
(21) $\text{Nullspace}(I^{2h}_{h})\;\bot\;\text{Range}(I^{h}_{2h}),$
(22) $\Omega^{h}=\text{Range}(I_{2h}^{h})\oplus\text{Nullspace}(I_{h}^{2h}A^{h}),$
(23) $\mathbf{e}^{h}=\mathbf{s}^{h}+\mathbf{t}^{h},$
Input : Grid level $\Omega^{h}$ with operator $A^{h}$; initial guess $v^{h}$; right-hand side $f^{h}$; pre-smoothing steps $\alpha_{1}$; post-smoothing steps $\alpha_{2}$.
Output : Updated approximation $v^{h}$.
if $\Omega^{h}$ is the coarsest grid then
   Relax $\alpha_{2}$ times on $A^{h}u^{h}=f^{h}$ with initial guess $v^{h}$;
end if
else
   Relax $\alpha_{1}$ times on $A^{h}u^{h}=f^{h}$ with initial guess $v^{h}$;
 $r^{h}\leftarrow f^{h}-A^{h}v^{h}$;
 // Compute the residual
 $f^{2h}\leftarrow I^{2h}_{h}r^{h}$;
 // Restrict the residual to the coarser grid
 $v^{2h}\leftarrow 0$;
 // Initial guess for the coarser grid
 $v^{2h}\leftarrow\text{Vcycle}(v^{2h},f^{2h})$;
 // Recursive call
 $v^{h}\leftarrow v^{h}+I^{h}_{2h}v^{2h}$;
 // Prolongate the correction and update
   Relax $\alpha_{2}$ times on $A^{h}u^{h}=f^{h}$ with initial guess $v^{h}$;
end if
return $v^{h}$;
Algorithm 1 V-cycle
Appendix B Multiwavelets
(a) The two-dimensional discrete wavelet transform.
(b) Matrix representation of decomposition.
Figure 5. (a) Two-dimensional discrete wavelet transform with $G$ (high-pass) and $H$ (low-pass) filters. Horizontal lines represent frequency bands; vertical lines show decomposition levels. (b) Matrix representation of decomposition, dividing $L_{n}$ into submatrices $B_{n}$, $C_{n}$, and $D_{n}$.
B.1. Multiresolution analysis
(24) $\mathbf{V}_{n}^{k}=\{$ $f:\text{the restriction of }f\text{ to the interval }(2^{-n}l,2^{-n}(l+1))$
is a polynomial
$\text{of degree less than }k\text{, for }l=0,\ldots,2^{n}-1$
, $\text{, and }f\text{ vanishes elsewhere}\}.$
The space $\mathbf{V}_{n}^{k}$ has dimension $2^{n}k$ and
(25) $\mathbf{V}_{0}^{k}\subset\mathbf{V}_{1}^{k}\subset\ldots\subset\mathbf{V}_{n}^{k}\subset\ldots.$
(26) $\mathbf{V}_{n}^{k}\oplus\mathbf{W}_{n}^{k}=\mathbf{V}_{n+1}^{k},\quad\mathbf{W}_{n}^{k}\bot\mathbf{V}_{n}^{k},$
and note that $\mathbf{W}_{n}^{k}$ is of dimension $2^{n}k$. Therefore, we have
(27) $\mathbf{V}_{n}^{k}=\mathbf{V}_{0}^{k}\oplus\mathbf{W}_{0}^{k}\oplus\mathbf{W}_{1}^{k}\oplus\ldots\oplus\mathbf{W}_{n-1}^{k}.$
(28) $\langle{f,g}\rangle=\int_{0}^{1}{f(x)g(x)dx}.$
(29) $\phi_{jl}^{n}=2^{n/2}\phi_{j}(2^{n}x-l),\quad j=0,\ldots,k-1,\quad l=0,\ldots,2^{n}-1.$
B.2. Multiwavelets
(30) $\int_{0}^{1}{\psi_{i}(x)\psi_{j}(x)dx}=\delta_{ij}.$
(31) $\int_{0}^{1}{\psi_{j}(x)x^{i}dx}=0,\quad i,j=0,1,\ldots,k-1.$
(32) $\psi_{jl}^{n}(x)=2^{n/2}\psi_{j}(2^{n}x-l),\quad j=0,\ldots,k-1,\quad l=0,\ldots,2^{n}-1,$
and supp $(\psi_{jl}^{n})=I_{nl}$ denotes the interval $2^{-n}l,2^{-n}(l+1)]$. The condition of orthonormality of $\psi_{0},\ldots,\psi_{k-1}$ yields
(33) $\int_{0}^{1}{\psi_{il}^{n}(x)\psi_{jm}^{n^{\prime}}(x)dx}=\delta_{ij}\delta_{lm}\delta_{nn^{\prime}}.$
B.3. Multiwavelet Transformation
(34) $\phi_{i}(x)=\sqrt{2}\sum_{j=0}^{k-1}{(h_{ij}^{(0)}\phi_{j}(2x)-h_{ij}^{(1)}\phi_{j}(2x-1))},\quad i=0,\ldots,k-1,$
(35) $\psi_{i}(x)=\sqrt{2}\sum_{j=0}^{k-1}{(g_{ij}^{(0)}\phi_{j}(2x)+g_{ij}^{(1)}\phi_{j}(2x-1))},\quad i=0,\ldots,k-1,$
where the coefficients $g_{ij}^{(0)}$, $g_{ij}^{(1)}$ depend on the choice of the order $k$. The function $\sqrt{2}\phi_{0}(2x)\ldots,\sqrt{2}\phi_{k-1}(2x)$ in (34) are orthonormal on the interval $0,\frac{1}{2}]$ whereas $\sqrt{2}\phi_{0}(2x-1)\ldots,\sqrt{2}\phi_{k-1}(2x-1)$ are orthonormal on the interval $\frac{1}{2},1]$. The matrices of coefficients
(36) $H^{(0)}=\{h_{ij}^{(0)}\},\quad H^{(1)}=\{h_{ij}^{(1)}\},\quad G^{(0)}=\{g_{ij}^{(0)}\},\quad G^{(1)}=\{g_{ij}^{(1)}\}$
By construction, we have $\langle{\phi_{i},\phi_{j}}\rangle=\delta_{ij}$, $\langle{\psi_{i},\psi_{j}}\rangle=\delta_{ij}$, and $\langle{\phi_{i},\phi_{j}}\rangle=0$ for $i,j=0,\ldots,k-1$, where $\langle,\rangle$ is the inner product defined in (28). Using this orthogonality conditions and (34-35), we obtain
(37) $H^{(0)}H^{(0)T}+H^{(1)}H^{(1)T}$ $=I,$
(38) $G^{(0)}G^{(0)T}+G^{(1)}G^{(1)T}$ $=I,$
(39) $H^{(0)}G^{(0)T}+H^{(1)}G^{(1)T}$ $=0.$
(40) $H$ $=\begin{pmatrix}H^{(0)}&H^{(1)}\end{pmatrix}$
$G$ $=\begin{pmatrix}G^{(0)}&G^{(1)}\end{pmatrix}$
(41) $H^{T}H+G^{T}G$ $=I$
$HG^{T}=GH^{T}$ $=0$
$HH^{T}=GG^{T}$ $=I$
(42) $H$ $=H_{y}\otimes H_{x}$
where $\otimes$ is the Kronecker product.
We denote the multiwavelet transform as
(43) $\mathcal{W}=\begin{pmatrix}H\\ G\end{pmatrix}$,
Due to the properties of $H$ and $G$ , $\mathcal{W}$ is orthogonal.
Appendix C Proof of Theorem 3.1
We show that the Multiwavelet–Based Multigrid Neural Operator (M2NO) can approximate any continuous operator $O^{*}:X\!\to\!Y$ to arbitrary accuracy, where $X=H^{s}(\Omega)$, $Y=H^{s^{\prime}}(\Omega)$ for $s,s^{\prime}\geq 1$, and $\Omega\subset\mathbb{R}^{d}$ is a bounded Lipschitz domain. Throughout, $\|\!\cdot\!\|_{X}$ and $\|\!\cdot\!\|_{Y}$ denote the respective Sobolev norms, and $L=\|O^{*}\|_{\mathcal{L}(X,Y)}$.
Step 1: Finite–Element Projection (C.1)
Because piecewise–polynomial spaces are dense in $H^{s}(\Omega)$, for every $\varepsilon>0$ there exists a mesh parameter $h>0$ and a finite–dimensional subspace $V_{h}=\operatorname{span}\{\psi_{1},\dots,\psi_{k}\}\subset X$ such that
(44) $\sup_{u\in C}\inf_{v_{h}\in V_{h}}\|u-v_{h}\|_{X}\;<\;\frac{\varepsilon}{3L},$
where $C\subset X$ is any compact set of interest (e.g. the training support). Let $\Pi_{h}\!:X\!\to\!V_{h}$ be the XX–orthogonal projection. Standard finite–element theory (13:fem) then yields
(45) $\|O^{*}(u)-O^{*}(\Pi_{h}u)\|_{Y}\;\leq\;L\|u-\Pi_{h}u\|_{X}\;<\;\frac{\varepsilon}{3}\quad\forall\,u\in C.$
Step 2: Operator Discretisation (C.2)
Because $O^{*}(V_{h})\subset Y$ lies in a finite–dimensional subspace, there exist basis functions $\{\phi_{1},\dots,\phi_{m}\}\subset Y$ and continuous coefficient maps $f_{i}\!:V_{h}\!\to\!\mathbb{R}$ such that, for every $v_{h}\in V_{h}$,
(46) $\Bigl\|O^{*}(v_{h})-\sum_{i=1}^{m}f_{i}(v_{h})\,\phi_{i}\Bigr\|_{Y}\;<\;\frac{\varepsilon}{3}.$
Inequality (46) follows from the Weierstrass approximation theorem applied on the compact set $O^{*}\bigl(C\bigr)$.
Step 3: M2NO Construction (C.3)
M2NO realises a map $\widehat{O}_{n}\!:V_{h}\!\to\!\operatorname{span}\{\phi_{i}\}_{i=1}^{m}$ by three modules: (i) projection $u\mapsto\Pi_{h}u$; (ii) multiwavelet restriction/prolongation, giving a multi–resolution representation of the coefficient vector $x=\bigl(\langle\Pi_{h}u,\psi_{j}\rangle\bigr)_{j=1}^{k}$; and (iii) a learnable linear map $A_{n}=[a_{ij}]_{i=1,\dots,m}^{j=1,\dots,k}$ trained to minimise
$E_{n}\;=\;\sup_{x\in\mathbb{R}^{k},\,\|x\|\leq 1}\Bigl\|\sum_{j=1}^{k}x_{j}\psi_{j}-\widehat{O}_{n}\!\!\left(\sum_{j=1}^{k}x_{j}\psi_{j}\right)\Bigr\|_{Y}.$
Universal approximation of linear maps by neural networks ensures that we can choose $n$ large enough so that $E_{n}<\varepsilon/3$. Equivalently,
(47) $\sup_{v_{h}\in V_{h},\,\|v_{h}\|_{X}\leq 1}\|O^{*}(v_{h})-\widehat{O}_{n}(v_{h})\|_{Y}\;<\;\frac{\varepsilon}{3}.$
Step 4: Wavelet Error Decomposition (Eq. 50)
Error residuals between $O^{*}$ and $\widehat{O}_{n}$ decompose into wavelet scales:
$O^{*}(u)-\widehat{O}_{n}(u)\;=\;r_{0}\;+\;\sum_{\ell=1}^{L}r_{\ell},$
[…]
(48) $\sup_{u\in C}\|O^{*}(u)-\widehat{O}_{n}(u)\|_{Y}\;\leq\;\frac{\varepsilon}{3}\;+\;\frac{\varepsilon}{3}\;+\;\frac{\varepsilon}{3}\;=\;\varepsilon.$
Conclusion.
For any $\varepsilon>0$, there exists a mesh size $h$, wavelet depth $L$, and network width $n$ such that (48) holds. Hence M2NO is a universal approximator for continuous operators $O^{*}\!:H^{s}(\Omega)\!\to\!H^{s^{\prime}}(\Omega)$.